Reliability of the Models and Their Match with Observations

Steven E. Brown^{∗} and Gerard M. Steiner^{†}

∗National Center for Atmospheric Research^{1} Boulder Colorado 80307 †National Standards Institute, Boulder, Colorado 11543

Abstract. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

INTRODUCTION

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system [1–3].

ΔT_{2c} = \frac{S_{C}}{S_{T}} ln(2)  (1)

∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

× (∑_{i<j}) ∑_{perm} \frac{1}{S_{12}} \frac{1}{S_{12}} ∑_{τ} c_{τ}^{f} .  (2)

1) The National Center for Atmospheric Research is sponsored by the National Science Foundation.

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

ΔT_{2c} = ∫_{0}^{t} F (τ) [\frac{S_{C}}{S_{T}} ] − S_{T} ΔT (t)  (3)

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

SPECIAL NUMBERING AND EXTRA SYMBOLS

To obtain numbers not normally produced by the automatic numbering, use the \eqnum{#1} command, where #1 is the desired equation number. For example, to get an equation number of (1.2′),

g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} ….  (1.2′)

A few notes on \eqnum{#1}. The \eqnum{#1} must come before the \label{#1}, if any. The numbering set with \eqnum{#1} is transparent to the automatic numbering in REVTEX; therefore, you must know the number ahead of time, and must make sure that the number set with \eqnum{#1} stays in step with the automatic numbering. \eqnum{#1} works with both single-line and multiline equations. You could, if you wished, do all the numbering in a paper manually with \eqnum{#1}.

Enclosing single-line and multiline equations in \begin{mathletters} and \end{mathletters} will produce a set of equations that are “numbered” with letters, as shown in Eqs. (4a) and (4b) below:

\{abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} ,  (4a)

M = ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}).  (4b)

If you use a \label{#1} command right after the \begin{mathletters}, then \ref{#1} can be used to reference all the equations in a mathletters environment. For example, the equations in the preceding mathletters environment were Eqs. (4).

The AIP book style will support all the “normal” TEX and L^{A}TEX symbols, and in addition, the following symbols

\corresponds  =^{△}  \gtrsim     > ∼
\lesssim         < ∼  \precsim ≺_{∼}
\succsim  ≻_{∼}  \lambdabar λ

\tensor{T}       T^{↔}

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system

POLICY-ANALYTIC APPROACH

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. See Table 1.

Atmospheric Model

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system [2].

Quantitative Differences

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

High-T_{c} calculations. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

TABLE 1. This table does not expand to the table does not ah ah ah expand to ta ble does NOT expand to textwidth.

Main sample  Secondary result  Previous result^{a}  Change^{b}(%)
Harkonen et al.  34.5  24.5  −23
Deckard and Batty  24.5  28.5  −13
Marcin  134.5  999999224.5  53

a First demonstrate that state-of-the-model climate models are credible simulators of the true climate system. b Estimated 2.

High-T_{c} calculations. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true dynamic climate system [3].

REFERENCES

1. Smith A., and Doe B., J. Chem. Phys. 76, 4056 (1982).

2. Jones, C., J. Chem. Phys. 68, 5298 (1978).

3. Jones, C., and Smith, A., Title of Book, New York: IEEE Press, 1978, ch. 6, pp. 23-26.

Reliability of the Models and Their Match with Observations

Steven E. Foo^{∗} and Gerard M. Bar^{†}

∗National Center for Atmospheric Research^{2} Boulder Colorado 80307 †National Standards Institute, Boulder, Colorado 11543

Abstract. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

INTRODUCTION

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system [1–3]. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

ΔT_{2c} = \frac{S_{C}}{S_{T}} ln(2)  (1)

∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

× (∑_{i<j}) ∑_{perm} \frac{1}{S_{12}} \frac{1}{S_{12}} ∑_{τ} c_{τ}^{f} .  (2)

2) The National Center for Atmospheric Research is sponsored by the National Science Foundation.

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

ΔT_{2c} = ∫_{0}^{t} F (τ) [\frac{S_{C}}{S_{T}} ] − S_{T} ΔT (t)  (3)

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system. See Figure 1.

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

SPECIAL NUMBERING AND EXTRA SYMBOLS

To obtain numbers not normally produced by the automatic numbering, use the \eqnum{#1} command, where #1 is the desired equation number. For example, to get an equation number of (1.2′),

FIGURE 1. Concentration profile climatology and climate model situation.

g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} ….  (1.2′)

A few notes on \eqnum{#1}. The \eqnum{#1} must come before the \label{#1}, if any. The numbering set with \eqnum{#1} is transparent to the automatic numbering in REVTEX; therefore, you must know the number ahead of time, and must make sure that the number set with \eqnum{#1} stays in step with the automatic numbering. \eqnum{#1} works with both single-line and multiline equations. You could, if you wished, do all the numbering in a paper manually with \eqnum{#1}.

Enclosing single-line and multiline equations in \begin{mathletters} and \end{mathletters} will produce a set of equations that are “numbered” with letters, as shown in Eqs. (4a) and (4b) below:

\{abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} ,  (4a)

M = ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}).  (4b)

If you use a \label{#1} command right after the \begin{mathletters}, then \ref{#1} can be used to reference all the equations in a mathletters environment. For example, the equations in the preceding mathletters environment were Eqs. (4).

The AIP book style will support all the “normal” TEX and L^{A}TEX symbols, and in addition, the following symbols

\corresponds  =^{△}  \gtrsim     > ∼
\lesssim         < ∼  \precsim ≺_{∼}
\succsim  ≻_{∼}  \lambdabar λ

\tensor{T}       T^{↔}

POLICY-ANALYTIC APPROACH

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. See Table 1.

Atmospheric Model

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system [2]. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system.

TABLE 1. This table does not expand to the table does not ah ah ah expand to ta ble does NOT expand to textwidth.

Main sample  Secondary result  Previous result^{a}  Change^{b}(%)
Harkonen et al.  34.5  24.5  −23
Deckard and Batty  24.5  28.5  −13
Marcin  134.5  999999224.5  53

a First demonstrate that state-of-the-model climate models are credible simulators of the true climate system. b Estimated 2.

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

Quantitative Differences

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

High-T_{c} calculations. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

High-T_{c} calculations. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true dynamic climate system [3].

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true dynamic climate system.

Qualitative Differences

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system.

REFERENCES

1. Smith A., and Doe B., J. Chem. Phys. 76, 4056 (1982).

2. Jones, C., J. Chem. Phys. 68, 5298 (1978).

3. Jones, C., and Smith, A., Title of Book, New York: IEEE Press, 1978, ch. 6, pp. 23-26.